OpenStudy (anonymous):

What is the solution set of x for the given equation?x^(3/2)-x^(1/3)+4=6

OpenStudy (anonymous):

-2, -1 2, -1 -8, -1 8, -1 2, 8

OpenStudy (anonymous):

\[x ^{3/2}-x ^{1/3}+4=6\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@a

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

I could never put ^1/3 in the equation on wolfram

OpenStudy (anonymous):

I plugged in 2,8 and got 6 somehow

OpenStudy (campbell_st):

are you sure the 1st term isn't \[x^{\frac{2}{3}} \] its just that it makes more sense that way...

OpenStudy (campbell_st):

because if the equation is \[x^{\frac{2}{3}} - x^{\frac{1}{3}} + 4 = 6\] it can be reduced to a quadratic... \[(x^{\frac{1}{3}})^2 - x^{\frac{1}{3}} - 2 = 0\]

OpenStudy (anonymous):

wow you came too late

OpenStudy (campbell_st):

well that's ok... 2, 8 is not a solution..

OpenStudy (anonymous):

yea It is not